OpenStudy (anonymous):

find the derivative of f(x)= x/sqrt(x^2 - 4) the answer is -4/(x^2-4)^3/2 but how?

OpenStudy (anonymous):

quotient rule plus chain rule

OpenStudy (anonymous):

i tried that

OpenStudy (anonymous):

i must be doing something wrong

OpenStudy (anonymous):

\[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\] with \[f(x)=x,f'(x)=1,g(x)=\sqrt{x^2-4},g'(x)=\frac{x}{\sqrt{x^2-4}}\]

OpenStudy (anonymous):

\[\frac{\sqrt{x^2-4}-x\frac{x}{\sqrt{x^2-4}}}{x^2-4}\] \[\frac{\sqrt{x^2-4}-\frac{x^2}{\sqrt{x^2-4}}}{x^2-4}\] now last step is to multiply top and bottom by \[\sqrt{x^2-4}\] and you will have your answer

OpenStudy (anonymous):

denominator will be \[(x^2-4)^{\frac{3}{2}}\] numerator will be \[x^2-4-x^2=-4\] and that is what you have above

OpenStudy (anonymous):

omg i got it thank you

OpenStudy (anonymous):

yw